(1) find the minimum value of the function f(x,y) = x^2 + 2y^2 on the circle x^2 + y^2 =1
(2) let the function f be defined by f(x,y) = 3x^2 y^4 / (x^4 + y^8) (x,y) not equal to (0,0) and 0 , (x,y) = (0,0). Show that f has directional derivatives in all directions at (0,0)
1
Expert's answer
2020-05-21T12:45:45-0400

(1). Make the "L=(x,y,l)=x^2+2y^2+l(x^2+y^2-1)" where "l" is Lagrange maltiplier.

"L'_x=2x+2lx,L'_y=4y+2ly." Solving system "L'_x=L'_y=0" and "x^2+y^2=1" then

"l=-1,y=0,x=1,x=-1"

"L''_{x^2}=2+2l,L''_{xy}=0,L''_{y^2}=4+2l" then "d^2L=(2+2l)dx^2+(4+2l)dy^2" .If "l=-1"

then "d^2L>0" hence "(1,0),(-1,0)" is points of minimum "f(1,0)=f(-1,0)=1"

(2). "\lim_{x,y\to0} \frac{3x^2y^4}{x^4+y^8}=\lim_{x\to0}\frac{3k}{k^2+1}=\frac{3k}{k^2+1}" where "k=\frac{x^2}{y^4}." The limit of the "f(x,y)" depends

on the value of "k" . Hence "f" has directional derivatives

in all directions at "(0,0)."